OpenStudy (anonymous):

sec(tan^-1(1)) = ? without using calculator

OpenStudy (jdoe0001):

\(\bf tan^{-1}(1)=\theta\implies tan(\theta)=1\qquad thus\qquad \theta=\square ?\)

OpenStudy (anonymous):

but I cannot use calculator!!

OpenStudy (phi):

There are some angles people memorize sin, cos and tan of: 0º, 30º, 45º, 60º and 90º

OpenStudy (jdoe0001):

well, use a Unit Circle, I assume you have one

OpenStudy (anonymous):

theta = 45

OpenStudy (anonymous):

now sec??